OpenStudy (lilah):

Which of the two functions below has the smallest minimum y-value? f(x)=4(x-6)^4+1 g(x)=2x^3+28 A. g(x) B. The extreme minimum y-value for f(x) and g(x) is negative infinity C. There is not enough information to determine D. f(x)

OpenStudy (lilah):

is it D? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

why do you think D?

OpenStudy (sshayer):

\[f(x)\ge 1\] \[-\infty<g(x)< \infty\]

OpenStudy (zzr0ck3r):

One has no minimum value, and the other one does. Do you know which one has a minimum value?

OpenStudy (lilah):

Too late i alreadu put D v.v

OpenStudy (lilah):

what was the answer though??

OpenStudy (lilah):

already*

OpenStudy (zzr0ck3r):

I disagree

jimthompson5910 (jim_thompson5910):

If you don't have a set domain, then the g(x) would win in terms of having the smallest value. Though technically g(x) doesn't have a min at all. It heads off to negative infinity

OpenStudy (zzr0ck3r):

A function with no set domain makes no sense and I am sure we are to assume it is the real numbers

jimthompson5910 (jim_thompson5910):

sorry that's what I meant. I meant if you don't have a set interval, and the domain is the set of all reals, then g(x) would have the smallest min, so to speak

OpenStudy (lilah):

ohh darn it, thanks for explaining guys